OpenStudy (ttop0816):

If |u| = 12, |v| = 8, and u · v = 48, then the angle between |u| and |v| is _____. 0° 30° 60° 90°

OpenStudy (johnweldon1993):

Well considering \[\large \vec{a} \cdot \vec{b} = |a||b|cos\theta \] we know that \[\large cos\theta = \frac{\vec{a} \cdot \vec{b}}{|a||b|}\]

OpenStudy (ttop0816):

90??

OpenStudy (johnweldon1993):

Not quite what I get so \[\large \theta = cos^{-1}\frac{48}{96}\] \[\large \theta = ?\]

OpenStudy (ttop0816):

OMG ITS 60!

OpenStudy (johnweldon1993):

There we go :)